### v1

# MagicPose4D: Crafting Articulated Models with Appearance and Motion Control

Abstract
With the success of 2D and 3D visual generative models, there is growing interest in generating 4D content. Existing methods primarily rely on text prompts to produce 4D content, but they often fall short of accurately defining complex or rare motions. To address this limitation, we propose MagicPose4D, a novel framework for refined control over both appearance and motion in 4D generation. Unlike traditional methods, MagicPose4D accepts monocular videos as motion prompts, enabling precise and customizable motion generation. MagicPose4D comprises two key modules: (i) Dual-Phase 4D Reconstruction Module which operates in two phases. The first phase focuses on capturing the model’s shape using accurate 2D supervision and less accurate but geometrically informative 3D pseudo-supervision without imposing skeleton constraints. The second phase refines the model using more accurate pseudo-3D supervision, obtained in the first phase and introduces kinematic chain-based skeleton constraints to ensure physical plausibility. Additionally, we propose a Global-local Chamfer loss that aligns the overall distribution of predicted mesh vertices with the supervision while maintaining part-level alignment without extra annotations. (ii) Cross-category Motion Transfer Module leverages the predictions from the 4D reconstruction module and uses a kinematic-chain-based skeleton to achieve cross-category motion transfer. It ensures smooth transitions between frames through dynamic rigidity, facilitating robust generalization without additional training. Through extensive experiments, we demonstrate that MagicPose4D significantly improves the accuracy and consistency of 4D content generation, outperforming existing methods in various benchmarks. Equal Contribution    Equal Advise
1 Introduction
The 4D generation task involves creating a temporal sequence of 3D models of moving objects. Given the difficulty of the general problem, recent approaches have made use of pre-trained models, using prompts to convey the user’s intentions. Recently, there has been a significant focus on the use of text/image-prompts to describe appearance and motion Ling et al. (2023); Ren et al. (2023); Zhao et al. (2023). The general pipeline of such existing consists of two stpdf: (i) acquiring static geometry through a 3D generation (e.g., a text/image-to-3D) model Liu et al. (2023), which generates 3D representation such as meshes/implicit fields according to text/image prompts, and (ii) obtaining motion information via a video generation model Blattmann et al. (2023a), which generates video according to text/image prompts. This approach has achieved impressive 4D content results.
However, challenges lie in facilitating users to freely and precisely specify the articulated motion of a non-rigid 3D object, and the generation faithfully reflecting the prompts, in terms of accurately capturing desired object appearance, geometry, and motion. The main issues with current methods are the following: (i) Temporal inconsistency in 3D geometry: Most current video generation models fail to ensure temporal consistency of 3D geometry. This involves additional complexity when objects are articulated because the 3D shape and movement must change mutually consistently during motion. For instance, while using the existing methods for 4D animal generation, we have observed unnatural and implausible configurations, such as the number of limbs of the object changing across different frames. (ii) Limited to common motions: Current video generation models perform well in generating simple, subtle, and common actions, such as walking or small motions like shaking, but they struggle to satisfactorily generate more complex motions, e.g., involving large or uncommon movements such as "King Kong dancing Hip-Hop" and "pig running like a rabbit" as shown in Fig.6. (iii) Text is inadequate for accurately describing the details of motion: As shown in Tab.4, most existing methods use text as prompts to describe the desired motion, with some efforts Bahmani et al. (2024) allowing trajectory (root body transformation) control. However, they are unable to specify detailed motions, e.g., by showing a real-world animal’s movement (say, via a monocular short video).
To address these issues, we propose MagicPose4D, which enables detailed control over both appearance and motion. Unlike existing methods, which rely mostly on text descriptions as motion prompts and for which it is difficult to convey complex or rare motions, our approach accepts monocular videos or dynamic mesh sequences as motion prompts. This allows for more precise supervision and faithful 4D generation. MagicPose4D introduces two key modules: Dual-Phase 4D Reconstruction Module and Cross-category Motion Transfer Module. The first module estimates a sequence of 3D models of the object while also simultaneously estimating motion parameters from the motion prompts, and the second module transfers this motion to the target object, which is generated by a 3D generation model controlled by appearance prompts.
Dual-Phase 4D Reconstruction Module: Given the complexity, especially with articulated objects, we divide this module into two phases. First, we use image descriptors (e.g., segments, flow) as 2D supervision. Additionally, we employ relatively less accurate but geometrically informative pseudo-3D models as geometric priors (using image-to-3D models) to provide 3D supervision for 3D reconstruction. This phase does not impose constraints emanating from the articulated nature of the object, e.g., captured in the underlying skeletal structure and plausible changes in it during motion, allowing each part to learn arbitrary rotations ($\mathbf{R}$) and translations ($\mathbf{t}$) while focusing on learning the model’s shape. In the second phase, we ensure the physical plausibility of the motion by enforcing kinematic chain-based skeletal movement constraints. Additionally, we propose a Global-local Chamfer loss, which ensures that the overall distribution of predicted mesh vertices aligns with the supervision while maintaining part-level alignment, without requiring additional annotations. Cross-category Motion Transfer Module: This module achieves cross-species motion transfer by mapping the skeleton of one species to another by establishing joint and limb correspondences through a kinematic-chain-based representation of the skeleton. Our motion transfer module is non-training-based. It helps improve generalization and prevents the poor performance seen in many existing methods when tested on data with significant gaps from the training set. Also, we leverage dynamic rigidity Zhang et al. (2024b) to guarantee smoothness between frames, unlike the existing approaches Song et al. (2021, 2023), which perform frame-independent pose transfer.
The following are the main contributions of this paper: (i) A Novel 4D Generation Framework: Our new framework leverages monocular videos as motion prompts, providing more accurate and more precisely specifiable 4D action generation. (ii) Skeleton Based 4D Representation: By using skeleton-mediated geometric and 3D prior, we achieve more accurate motion estimation and 3D reconstruction, improving the physical plausibility of the generation. (iii) Global-local Chamfer Loss: We introduce a novel loss function to better align estimated mesh vertices with the supervisory 3D model overall while maintaining part-level alignment, without additional annotations. (iv) Cross-category Motion Transfer: Our cross-category mapping in terms of skeleton based dynamic rigidity representation enables smooth transitions between frames and robust generalization without additional training. (v) Outperforms SOTA: Through extensive experiments, we demonstrate that MagicPose4D provides highly accurate 4D content and significantly outperforms existing methods for 4D reconstruction and pose transfer across all three benchmarks that we experiment with.
Figure 1: Overview of MagicPose4D, which takes motion prompts (monocular video or dynamic mesh sequence) and appearance prompts (text or image) to control the 4D content generation.
2 Related Work & Motivation
4D/3D Reconstruction. Significant advancements in 4D/3D reconstruction include the development of specialized parametric models such as SCAPE Allen et al. (2003) and SMPL Loper et al. (2023a) for human bodies, MANO Romero et al. (2022) for hand movements, FLAME Li et al. (2017) and EMOCA Daněček et al. (2022) for facial expressions, and SMAL Zuffi et al. (2017) for quadruped animals. However, these methods require predefined parameters that include skeletons and skinning weights, limiting their generalization to uncommon species and out-of-distribution data. Recent neural implicit-based methods Palafox et al. (2021); Yang et al. (2022b); Giebenhain et al. (2023); Zhang et al. (2024b); Yang et al. (2021b) offer promising alternatives by jointly learning static 3D meshes and time-varying parameters without predefined templates. However, they often fail with monocular videos containing sparse views, limiting their practical applicability. To address this, MagicPose4D introduces an innovative Dual-Phase 4D Reconstruction module, which reduces the requirements from multi-view videos to single-view videos and achieves robust reconstruction.
Diffusion-based 4D Generation and Pose Transfer. Recent diffusion-based 4D generation methods have shown promising results by leveraging text-to-3D models followed by text-to-video supervision. These techniques Singer et al. (2023); Ren et al. (2023); Ling et al. (2023); Bahmani et al. (2023); Zheng et al. (2023); Gao et al. (2024); Yang et al. (2024); Jiang et al. (2023); Zhao et al. (2023) have improved the geometry and appearance of generated models but are generally limited to minor movements within a fixed location and rely on text prompts, making precise motion control difficult. Our goal is to create dynamic 4D animations that closely transfer the motion of any given reference real-world object. For pose transfer, recent skeleton-based frameworks Song et al. (2021); Liao et al. (2022) have explored the use of rigging points and key points. However, learning-based methods often struggle with in-the-wild data and significant domain gaps between identities. MagicPose4D introduces a cross-category motion transfer module that supports cross-species transfer while ensuring generalization and maintaining temporal smoothness.
3 MagicPose4D
MagicPose4D accepts two distinct types of input prompts: (i) appearance prompts and (ii) motion prompts. Consistent with recent methods Blattmann et al. (2023b); Tang et al. (2023), both images and textual descriptions can function as appearance prompts, delineating the desired object and its visual characteristics. In a departure from existing approaches, MagicPose4D enables users to specify precise motions and trajectories by providing a video/mesh sequence that represents the anticipated movement.
As illustrated in Fig 1, MagicPose4D comprises three critical components: (i) the 4D Reconstruction module (Sec.3.2), (ii) the Cross-Category Motion Transfer module (Sec.3.3), and (iii) the Image-to-3D Generation module. Each module is tailored to facilitate distinct aspects of dynamic modeling, enabling adaptive 4D reconstructions that align with user-defined specifications.
3.1 Terminology and Overview
To represent an animated 3D model, our method learns static representations, such as the visible canonical shape $\mathbf{S}\in\mathbb{R}^{N\times 3}$, the underlying skeleton $\mathbf{S_{k}}=\{\mathbf{J}\in\mathbb{R}^{J\times 3},\mathbf{B_{s}}\in\mathbb{R}^{B},\mathbf{P_{i}}\in\mathbb{R}^{B}\}$, and the skinning weights $\mathbf{W}\in\mathbb{R}^{N\times B}$. Additionally, it captures time-varying parameters, including the global-local (root body-bone) transformations $\mathbf{\tau}=\{\tau_{0}^{t},\tau_{1}^{t},\ldots,\tau_{B}^{t}\}$ and the camera parameters $\mathbf{P_{c}}$. Here, $B$, $N$, and $J$ represent the number of bones, vertices on the mesh surface, and joints, respectively. The transformations $\tau_{i}^{t}\in SE(3)$ include $\tau_{0}^{t}$ for the root body, with the remaining $\tau_{i}^{t}$ for the bones. The skeleton topology is described by $\mathbf{J}$ (joint coordinates), $\mathbf{B_{s}}$ (bone scale), and $\mathbf{P_{i}}$ (parent indices of joints).
As illustrated in Fig 1, utilizing a short monocular video as a motion prompt, we introduce a dual-phase 4D reconstruction module (Sec.3.2) designed to predict a sequence of 4D meshes as motion references. Appearance prompts may consist of text descriptions, images, or a monocular video. Depending on the type of input, a corresponding module—text-to-3D, image-to-3D, or dual-phase 4D reconstruction is employed to generate the static 3D representation of the desired object. Subsequently, this static representation along with the motion reference are fed into the cross-category motion transfer module (Sec.3.3). This module adeptly transfers motions from the reference to the target object while ensuring temporal smoothness and consistency.
Figure 2: Overview of Dual-Phase 4D Reconstruction Module.
3.2 Dual-Phase 4D Reconstruction from video
Reconstructing 4D models from short monocular videos is a challenging task as it requires jointly learning numerous parameters, resulting in an extensive optimization space. Erroneous predictions of any parameter can trigger cascading failures. To address these challenges, we propose a Dual-Phase 4D Reconstruction module, which employs differentiated supervision across two distinct phases and focuses on learning diverse representations.
In the first phase, the primary focus is on accurately capturing the external appearance (shape) of the model. The underlying skeleton serves merely as an intermediary variable, utilizing non-kinematic chains skeleton and learnable skinning weights to afford the skeleton greater deformation freedom. This approach accelerates the learning of correct shapes. In contrast, the second phase aims to derive a more physically plausible motion reference for effective motion transfer. Therefore, we adopt kinematic chain skeletons and heat diffusion-based skinning weights, which narrow the deformation space of the skeleton, thereby ensuring the plausibility of the internal structure. (Sec.3.2.1)
From a supervision perspective, the first phase blends 2D and pseudo-3D supervision, updating the pseudo-3D supervision at the end of this phase. In the second phase, the 2D loss is removed, relying solely on the updated 3D pseudo-supervision to guide the learning process. (Sec.3.2.2)
3.2.1 Model Articulation
Skinning Weights $\mathbf{W}$ is designed to represent the probabilities that each vertex corresponds to $B$ semi-rigid parts. In the first phase, following Yang et al. (2021b, 2022b), the skinning weights are modeled by the mixture of $B$ Gaussian ellipsoids as: $\mathbf{W}_{n,b}=\mathcal{F}e^{-\frac{1}{2}(\mathbf{X}_{n}-\mathbf{C}_{b})^{T}\mathbf{Q}_{b}(\mathbf{X}_{n}-\mathbf{C}_{b})},$ where $\mathcal{F}$ is the factor of normalization and precision matrices $\mathbf{Q}_{b}=\mathbf{V}_{b}^{T}\mathbf{\Lambda}_{b}\mathbf{V}_{b}$. In each Gaussian ellipsoid, $\mathbf{C}\in\mathbb{R}^{B\times 3}$ denotes Gaussian centers, $\mathbf{V}\in\mathbb{R}^{B\times 3\times 3}$ defines the orientation and $\mathbf{\Lambda}\in\mathbb{R}^{B\times 3\times 3}$ denotes the diagonal scale matrix. $\mathbf{X}_{n}$ is the 3D location of vertex n. In the second phase, as shown in Fig.3, a skeletonization module, is leveraged to obtain a skeleton for the canonical mesh, and following Tong et al. (2012); Baran and Popović (2007), skinning weights $\mathbf{W}$ are obtained by a heat diffusion process. This approach guarantees a more natural assignment of skinning weights to the bones, enhancing the realism of the skeletal animations.
Blend Skinning. The mapping from surface vertex $\mathbf{X}_{n}^{0}$ in canonical space (time 0 by default) to $\mathbf{X}_{n}^{t}$ at time t in camera space is designed by blend skinning. The forward blend skinning is shown below:
$\begin{split}\mathbf{X}_{n}^{t}=\mathbf{\tau}_{0}^{t}(\sum_{b=1}^{B}\mathbf{W}_{n,b}\mathbf{\tau}_{b}^{t})\mathbf{X}_{n}^{0},\end{split}\vspace{-0pt}$ (1)
including the number of bones $B$, skinning weights $\mathbf{W}$ and the transformation $\mathbf{\tau}^{t}=\{\mathbf{\tau}_{b}^{t}\}_{b=0}^{B}$ at time $t$, where $\mathbf{\tau}_{0}^{t}$ represents the root body transformation and $\mathbf{\tau}_{b>0}^{t}$ describes the bone transformation. Each $\mathbf{X}_{n}^{0}$ is first transformed by the weighted sum of each bone transformation $\mathbf{T}_{b>0}^{t}$ and then transformed by root body transformation $\mathbf{T}_{0}^{t}$ to achieve $\mathbf{X}_{n}^{t}$.
Skeleton Articulation. As described above, bone and root body transformations $\mathbf{\tau}$ are crucial in the process of blending skinning. In the first phase, we define these transformations as independently learnable $SE(3)$ transformations. Practically, this is implemented by initializing learnable quaternions to compute the rotation matrices, along with defining learnable translations to achieve affine transformations. During this phase, we impose no constraints on the skeleton; each bone can rotate and translate freely. These bones serve merely as intermediate variables, with the ultimate objective of accurately determining the correct shape. In the second phase, we define per-frame joint angles $Q$, each describing the pose between a bone and its parent with three degrees of freedom. Instead of directly learning bone transformations, we determine $Q$ and compute the bone transformations $\mathbf{\tau}_{b>0}^{t}$ using forward kinematics. The deformed mesh is then obtained by blend skinning.
Extendable Bones. It is noteworthy that kinematic-chain-skeleton-based methods rest on a fundamental assumption that the structure of the utilized skeleton closely mirrors the natural skeletal architecture of animals. This allows for the desired deformed mesh to be achieved by manipulating the skeleton’s pose and employing blend skinning techniques. However, in practice, this assumption often does not hold completely. Animal bones are not solely rigid hinge connections; they include extendable cartilaginous tissues that lead to non-rigid deformations among joints. To address this, we introduce the concept of extendable bones. We allow for slight variations in bone length between frames, achieved by learning a time-varying scale parameter for each bone $\mathbf{B_{s}}\in\mathbb{R}^{B}$. This modification enhances the flexibility and realism of our skeletal model.
3.2.2 Supervision and Losses
As shown in Fig. 2, given the input from a monocular video, we utilize segmentation models Kirillov et al. (2023); Zhang et al. (2024a) and optical flow prediction models Teed and Deng (2020) to obtain 2D supervision. Furthermore, we employ an image-to-3D model to independently predict meshes for each frame, serving as pseudo-3D supervision. A question arises: "Why not directly use an image-to-3D model to independently predict 3D meshes for each frame?" As illustrated in Fig.7 (c), objects always exhibit self-occlusion in the video, under which circumstances image-to-3D models typically fail to produce accurate results. For instance, some frames might only depict a camel with two visible legs, resulting in a 3D mesh sequence that does not effectively capture the action information portrayed in the video. Moreover, as demonstrated in Video 3 of the supplementary materials, meshes generated directly from the image-to-3D module are independent of one another, thus lacking temporal continuity and smoothness. Although the initial pseudo-3D supervision may not be very accurate, it still provides valuable geometric priors that help address the information loss caused by insufficient perspectives of the object in the video.
In the first phase, we blend both 2D and 3D supervision to optimize the mesh shape and leverage a reconstruction loss, which consists of silhouette loss, optical flow loss, texture loss, perceptual loss, smooth, motion, and symmetric regularizations, and global-local chamfer (GLC) Loss. In the second phase, we only leverage GLC Loss and regularization terms without 2D losses. The details of the 2D loss functions and regularizations will be described in the Appendix.6.5.
Global-Local Chamfer Loss (GLC). The objective of GLC loss is to ensure that the predicted mesh closely resembles the expected mesh in (i) overall shape and in terms of their (ii) respective poses. This dual focus helps achieve high fidelity in both the structural and positional accuracy of the meshes. The GLC loss is the sum of chamfer distances across two levels. Initially, it involves the computation of the chamfer distance for the entire predicted mesh $\mathbf{S}$ and Pseudo mesh $\mathbf{\hat{S}}$ as follows:
$\begin{split}\mathcal{L}_{\text{global}}(\mathbf{S},\mathbf{\hat{S}})=\frac{1}{|\mathbf{S}|}\sum_{x\in\mathbf{S}}\min_{y\in\mathbf{\hat{S}}}\|x-y\|^{2}+\frac{1}{|\mathbf{\hat{S}}|}\sum_{y\in\mathbf{\hat{S}}}\min_{x\in\mathbf{S}}\|x-y\|^{2}.\end{split}$ (2)
The second level involves computing the weighted chamfer distances between $B$ parts. First, we calculate skinning weights $\mathbf{W},\mathbf{\hat{W}}\in\mathbb{R}^{N\times B}$ for $\mathbf{S}$ and $\mathbf{\hat{S}}$ via the heat diffusion process Tong et al. (2012); Baran and Popović (2007). Then, we perform an argmax across the $B$ dimension to achieve a part-wise decomposition of the whole body into $K$ parts, where $K$ is less than or equal to $B$. In practice, $K$ often equals $B$, but in cases where $K$ is less than $B$, the loss computation simply omits the non-existent parts. Subsequently, we calculate the part-level Chamfer distances between the $K$ pairs from predicted mesh $\mathbf{P}_{k},k\in\{0,...,K-1\}$ and Pseudo mesh $\mathbf{\hat{P}}_{k}$, as shown in the following equation:
$\begin{split}\mathcal{L}_{\text{local}}(\mathbf{S},\mathbf{\hat{S}})=\frac{1}{|K|}\sum_{k=0}^{K-1}\frac{1}{|\mathbf{P}_{k}|}\sum_{x\in\mathbf{P}_{k}}\mathcal{W}\min_{y\in\mathbf{\hat{P}}_{k}}\|x-y\|^{2}+\frac{1}{|\mathbf{\hat{P}}_{k}|}\sum_{y\in\mathbf{\hat{P}_{k}}}\mathcal{W}\min_{x\in\mathbf{P}_{k}}\|x-y\|^{2},\end{split}$ (3)
where $\mathcal{W}=\mathbf{W}_{x,k}\times\mathbf{\hat{W}}_{y,k}$ represents the multiplication of skinning weights between vertices $x$ and $y$ for bone $k$. $\mathcal{W}$ serves as a rough key points estimator based on skinning confidence. It assigns greater weight to vertices at the central part of a segment and lesser weight to vertices at the junctions of multiple parts. This approach effectively addresses inconsistencies at the edges of part decompositions when calculating skinning weights using heat diffusion for models in various poses.
Figure 3: Overview of Cross-Category Motion Transfer Module.
3.3 Cross-Category Motion Transfer
In the second phase of 4D Reconstruction, as shown in Fig. 3(a), we extract the underlying skeletal motion of the reference meshes. Given a sequence of meshes for the reference object, the skeletonization module extracts the skeleton of the canonical shape, which is defined as the shape corresponding to time $=0$. We fix the canonical skeleton and the skinning weights, thus controlling the skeleton’s pose solely by learning the pre-frame angles and bone scales. Subsequently, by employing blend skinning, we obtain the deformed mesh.
The input to the skeletonization module is a mesh, along with an optional skeleton template. When the reference object is a commonly recognized form, such as a quadruped or a human, we utilize the corresponding skeleton template and embed it into the mesh. Following the methodology described in Baran and Popović (2007), we construct a coarse discrete representation to locate the approximate position of the skeleton within the internal space of the mesh. This process involves embedding the skeleton into a graph derived from the character’s internal volume and determining an optimal solution using the A* algorithm across all possible matches. When a template of the object is not available, we employ skeleton extraction methods Bærentzen and Rotenberg (2021); Zhang et al. (2024b) to extract the canonical skeleton of the reference object. After learning the reference motion represented by the pre-frame angles and the bone scale of the skeleton, we can readily compute the global-local transformation of the target object using forward kinematics, as illustrated in Fig. 3(b). We extract the canonical skeleton for the target object by inputting the canonical skeleton of the reference object, adhering to the method previously described following Baran and Popović (2007). Subsequently, blend skinning is employed to generate the deformed meshes.
4 Experiments
This section primarily covers the following parts: (i) 4D reconstruction results and comparisons; (ii) motion transfer results and comparisons; and (iii) our framework v.s. 4D generation. Detailed descriptions of the benchmarks (Sec.6.2), implementation (Sec.6.3), and as shown in Tab.5, more than $60$ video results (Sec.6.1) are provided in the Appendix.
Figure 4: Appearance and Motion Controlled 4D Generation.
4.1 Qualitative and Quantitative Comparison
4D Reconstruction. As depicted in Fig.5, we first compare the results of MagicPose4D with LASR Yang et al. (2021a) and BANMo Yang et al. (2022a). These methods fall short of achieving ideal reconstruction, due to a lack of structural information about the objects, since they learn from only a single monocular video, which contains sparse views. NeRF-based methods such as BANMo Yang et al. (2022b) and MagicPony Wu et al. (2023) require particularly dense views from multiple cameras or long videos to achieve decent results. However, MagicPose4D achieves good reconstructions from only sparse views.
Furthermore, for a comprehensive quantitative comparison, we examine the performance of MagicPose4D against the best existing methods, LASR Yang et al. (2021a) and ViSER Yang et al. (2021c), which perform best (in terms of 2D keypoint transfer accuracy) among existing methods using a single monocular video as input. As shown in Tab.1, MagicPose4D consistently outperforms both LASR and ViSER for all animal subjects with a notable margin. When we run the author-provided codes for LASR, and ViSER on DAVIS, and PlanetZoo, we find that their results are highly variable across different runs. The results we report for these cases are therefore the mean accuracies we have obtained over multiple runs of all three methods.
Figure 5: 4D Reconstruction Results. We show the mesh reconstruction results of (a) LASR, (b) BANMo, and (c) Ours in the PlanetZoo’s bear, zebra, elephant, and giraffe.
Figure 6: 4D Generation Comparison of MagicPose4D with Animate124 Zhao et al. (2023), Motion Transfer Comparison of MagicPose4D with 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023). Videos are in Sec.6.1.
Cross-Category Motion Transfer. We present motion transfer results from MagicPose4D in Fig. 4. The target identities and reference motion sequences can either be humanoid or animal subjects. MagicPose4D is able to learn and retarget the motion from dynamic mesh prompts, e.g., Hip-Hop Dancing, or monocular video prompts, e.g., Hands Up.
We further compare the motion transfer ability for mesh generation quality with recent methods 3D-CoreNet Song et al. (2021) and X-DualNet Song et al. (2023) in Fig. 6 and Tab.2 (c). Both methods use a deep neural network to learn latent shape codes to retarget the motions and are trained/evaluated on SMPL Loper et al. (2023b)(humanoid) and SMAL Zuffi et al. (2017)(animal). Since they do not include disentangled components or shape deformation modules to represent the structural information of 3D meshes, these methods rely heavily on pre-processed training data with high-quality mesh annotations. Due to these constraints, these baselines cannot generalize well to in-the-wild target identities or reference motions. In contrast, MagicPose4D generates temporally consistent and smooth motions, while strictly preserving the identity and appearance of the target mesh. It is also worth noting that because previous methods focus only on pose transfer without considering the motion trajectory, the generated object always stays at the same position. This can be observed in side-by-side video comparisons in the supplementary materials Sec.6.1.
Our Framework v.s. Existing 4D Generation. Text/image-prompts-based 4D generation has been a popular trend. To compare with a representative recent work Animate124 Zhao et al. (2023), we feed the reference image, which provides the identity, and text prompt, which describes the motion of the generated object, into their model and optimize the NeRF with SDS-loss. In our case, we use the reference image to generate the target mesh with an Image-to-3D model and transfer the motion of the reference mesh sequence. We visualize the rendered video from different viewpoints from Animate124 and compare it with MagicPose4D in Fig. 6(a) and Fig. 6(b). More video comparisons are in the supplementary material. MagicPose4D provides more temporally consistent and smooth generation with the learning of motion deformation. Since there are no ground truth mesh sequences to evaluate the generation, we provide a comprehensive user study for Motion Transfering and 4D Generation for a qualitative evaluation in Sec. 6.4 and Tab.3 to conclude our findings.
Table 1: 2D Keypoint Transfer Accuracy. This table presents the 2D keypoint transfer accuracy on DAVIS and PlanetZoo datasets. For classes from DAVIS (camel) and PlanetZoo (giraffe, tiger), we carry out multiple executions and report the mean of the observed accuracy in each case. Note that the results for LASR and ViSER here are from our executions of the codes provided by the authors.
Method DAVIS PlanetZoo
camel cow dog bear dance-twirl giraffe tiger elephant bear zebra deer Ave.
ViSER 71.7 73.7 65.1 72.7 78.3 51.2 68.4 68.9 60.3 55.7 57.3 65.8
LASR 75.2 80.3 60.3 83.1 55.3 56.3 70.4 69.5 63.1 57.4 60.3 66.5
MagicPose4D 78.9 83.1 67.9 86.2 84.3 62.9 73.3 70.6 67.1 60.2 61.9 72.4
Figure 7: Ablation Experiments.
Table 2: User study of MagicPose4D on Motion Transfer. We collect the rating results from 50 participants of eleven mesh-sequence comparison experiments. The scale of rating is 0 (low) - 5 (high). The participants found that MagicPose4D generates the best motion transfer results. Criteria for judgment: 1) The generated motion should match the reference pose mesh sequence. 2) The identity of the transferred mesh should match the identity reference. 3) The generated mesh sequence should be consistent and smooth. More details can be found in Appendix Sec.6.4
Method Exp-1 Exp-2 Exp-3 Exp-4 Exp-5 Exp-6 Exp-7 Exp-8 Exp-9 Exp-10 Exp-11 Ave.
3D-CoreNet Song et al. (2021) 1.85 1.73 2.10 2.31 1.90 1.73 2.29 2.73 2.65 1.88 1.79 2.09
X-DualNet Song et al. (2023) 2.88 1.94 1.53 1.88 2.22 2.18 2.14 2.86 2.71 2.22 2.31 2.26
MagicPose4D 4.92 4.47 4.69 4.39 4.45 4.45 3.90 4.12 4.02 4.10 4.20 4.34
Table 3: User study of MagicPose4D on 4D Generation compared to Animate124 Zhao et al. (2023). The settings are the same as the study on Motion Transfer. Criteria for judgment: 1) The generation’s identity should match the reference image. 2) The video should be temporal-consistent and smooth.
Method Exp-1 Exp-2 Exp-3 Exp-4 Exp-5 Exp-6 Ave.
Animate124 Zhao et al. (2023) 2.60 1.98 1.90 1.92 2.28 2.24 2.15
MagicPose4D 4.53 4.37 3.91 4.00 4.60 4.62 4.34
4.2 Diagnostic
Effectiveness of Global-Local Chamfer Loss. Relying solely on the geometric information contained in a short monocular video often makes it challenging to achieve satisfactory 3D reconstruction results. This difficulty arises because most videos typically offer sparse viewpoints and the presence of non-rigid object deformations increases the optimization space. Consequently, many methods that rely only on 2D supervision Yang et al. (2021b, 2022b); Wu et al. (2023) struggle to ensure effectiveness on in-the-wild videos. To enhance the generalization capability of the reconstruction module, we utilize existing image-to-3D model methods to obtain geometric priors. By using these models, we generate pseudo-3D meshes for each frame as 3D supervision. We employ the Chamfer distance as the 3D loss objective to align the predicted 3D mesh closely with the pseudo-3D mesh in terms of point distribution. However, global Chamfer distance alone can only ensure similar point distributions between the two meshes and does not guarantee the correspondence of points as expected. For instance, as shown in Fig.7(a), without using the local Chamfer loss, the overall point distribution might be similar, but the mesh shape could be entirely incorrect, such as the left and right legs being swapped or vertices originally on the right front leg being moved near the left hind leg. By incorporating the local Chamfer loss, we enforce consistency between each part of the predicted mesh and the pseudo mesh, thus ensuring pose consistency and significantly mitigating the previously mentioned issues.
Temporal Consistency is crucial for 4D generation. As shown in Fig.7 (b), performing pose transfer independently for each frame results in noticeable discontinuities when viewing the entire video. To address this, we first define the canonical space using the initial frame, where all frames share the same canonical mesh, skeleton, and skinning weights. Each subsequent frame is derived from the deformation of the first frame using global-local transformations and blending skinning. This approach ensures the consistency of the static model and rigging. Additionally, we introduce a dynamic rigidity regularization term between consecutive frames, which minimizes the deformation of the object between successive frames. This ensures temporal smoothness of the 4D content.
5 Conclusion & Limitations
We introduce MagicPose4D, a novel framework for 4D generation providing more accurate and customizable 4D motion transfer. We propose a dual-phase reconstruction process that initially uses accurate 2D and pseudo 3D supervision without skeleton constraints and subsequently refines the model with skeleton constraints to ensure physical plausibility. We incorporate a novel loss function that aligns the overall distribution of mesh vertices with the supervision and maintains part-level alignment without additional annotations. MagicPose4D enables cross-category motion transfer using a kinematic-chain-based skeleton, ensuring smooth transitions between frames through dynamic rigidity and achieving robust generalization without the need for additional training.
The main limitations of our method and existing works are: (i) Deformation based on blend skinning relies on accurate and robust skeletons and skinning weights predictions, facing a trade-off between generalization and accuracy. Learning-based methods have limited generalization due to restricted training datasets, while non-learning methods suffer from inductive bias, leading to suboptimal results. (ii) MagicPose4D can infer poses quickly for pose transfer without training, but 4D reconstruction requires significant training (10 hours on an L40S). (iii) Our method struggles with detailed motion control, such as fingers and facial features, due to the challenge of capturing fine-grain details during 4D reconstruction. These issues represent future research directions.